The moon is observed from two diametrically opposite points A and B on earth . The angle  thita subtended at the moon by the two directions of observation is 1° 54' . Given the diameter of the earth to be about 1.276 × 10^7 m , compute the distance of the moon from the earth?

Grade:11

2 Answers

Khanshama
6 years ago
1 degree 54 minute(60+54)min=114 min`Now 114×60=6840"1"=4.85×10*-6radSo 6840×4.85×10*-6=3.32×10*-2Now D=diameter of earth /angle of parallax10276×10*7/3.32×10*-2=3.843×10*8m
max
5 years ago
1 degree 54 minute(60+54)min=114 min`Now 114×60=6840"1"=4.85×10*-6radSo 6840×4.85×10*-6=3.32×10*-2Now D=diameter of earth /angle of parallax10276×10*7/3.32×10*-2=3.843×10*8m n